OpenStudy (mew55):

Linear Algebra: Write down the 2X3 matrices A and B such that aij=i+j and bij=(-1)^i+j

jimthompson5910 (jim_thompson5910):

Hint: first row, first column of matrix A: a11 = 1+1 = 2 So the first row, first column element of matrix A is 2 First row, second column is a12 = 1+2 = 3 and so on